OpenStudy (asylum15):

Prove that the following integral = 0

OpenStudy (asylum15):

\[\frac{ 1 }{ 2\pi } \int\limits_{-\pi}^{\pi} f(t) dt\] where f(t) = 1

OpenStudy (deaconvex115):

now, seeing this problem, i feel,(dramatic pause)

OpenStudy (deaconvex115):

INCREDIBLY STUPID

OpenStudy (asylum15):

What?

OpenStudy (deaconvex115):

its compicated for me

OpenStudy (deaconvex115):

the only answer i can come up with is

OpenStudy (karim728):